# On computability of equilibrium states

### v2

@@ 4.1. Ruelle–Perron–Frobenius operators and cones
Note that $\overline{\varphi}(x)=\varphi(x)-P(T,\varphi)$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$. So for each $\overline{x}\in T^{-1}(x)$,
(4.8) $\exp(\overline{\varphi}(\overline{x}))\leqslant\exp(a_{0}\rho(\overline{x},y(\overline{x}))^{v_{0}})\cdot\exp(\overline{\varphi}(y(\overline{x})))\leqslant\exp(a_{0}\lambda^{-v_{0}}\rho(x,y)^{v_{0}})\cdot\exp(\overline{\varphi}(y(\overline{x}))).$
Moreover, by (2.16), for each $\overline{x}\in T^{-1}(x)$,
@@
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$, where the last equality holds due to the definitions of $\lambda_{1}$ and $a^{\prime}$. By induction, it follows from $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})\subseteq\mathcal{C}(a^{\prime},v_{0},b)$ that $L^{m}(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$. Finally, we establish that $L^{m}(u)$ satisfies property (iii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},v_{0},b^{\prime})$, namely,
@@
Without loss of generality, assume that $\inf_{x\in X}u(x)=1$. Let $x_{*}$ and $x^{*}$ be two points in $X$ satisfying that $u(x_{*})=1$ and $u(x^{*})=\sup_{x\in X}u(x)$. Since $u\in\mathcal{C}(a^{\prime},v_{0},b)$, we have $u(x^{*})\leqslant b$. First consider the case where $\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}\leqslant u(x^{*})\leqslant b$. Consider a point $w\in X$. Since $G^{\prime}$ is an $(\epsilon/2)$-net of $X$, there exists an integer $i\in[1,n^{\prime}]$ with $x_{*}\in B(x^{\prime}_{i},\epsilon/2)$. Since $\bigcap_{i=1}^{n^{\prime}}T^{m}(B(x^{\prime}_{i},\epsilon/2))=X$, there exists $\overline{w}_{*}\in T^{-m}(w)$ such that $\rho(\overline{w}_{*},x^{\prime}_{i})<\epsilon/2$. Hence, $\rho(\overline{w}_{*},x_{*})\leqslant\rho(\overline{w}_{*},x^{\prime}_{i})+\rho(x^{\prime}_{i},x_{*})<\epsilon\leqslant\xi$. Moreover, since $u\in\mathcal{C}(a^{\prime},v_{0},b)$, we have
@@
Hence, by the definitions of $\epsilon,\,Q$, $L(\mathbbm{1})=\mathbbm{1}$, (2.17), (4.5), and (4.11), we have
$L^{m}(u)(w)$ $=\frac{u(\overline{w}_{*})u_{\varphi}(\overline{w}_{*})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w}_{*})}+\sum_{\overline{w}\in T^{-m}(w)\setminus\{\overline{w}_{*}\}}\frac{u(\overline{w})u_{\varphi}(\overline{w})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w})}$
$\leqslant\frac{\exp(a^{\prime}\epsilon^{v_{0}})\cdot u_{\varphi}(\overline{w}_{*})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w}_{*})}+u(x^{*})\sum_{\overline{w}\in T^{-m}(w)\setminus\{\overline{w}_{*}\}}\frac{u_{\varphi}(\overline{w})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w})}$
@@
$\leqslant b-\biggl{(}\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}-\frac{b}{2}\biggr{)}C^{-2}Q^{m}$
$=\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2},$
where the last equality comes from direct calculation. Next, we consider the case where $u(x^{*})\leqslant\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}$. Since $L(\mathbbm{1})=\mathbbm{1}$, we have
$\sup_{w\in X}L^{m}(u)(w)\leqslant\sup_{w\in X}u(w)\leqslant\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}.$
To sum up, we always have $\sup_{w\in X}L^{m}(u)(w)\leqslant\frac{(C^{-2}Q^{m}+2)b}{2C^{-2}Q^{m}+2}$. Moreover, since $L(\mathbbm{1})=\mathbbm{1}$, it follows that $\inf_{w\in X}L^{m}(u)(w)\geqslant 1$. Therefore, we have estalished (4.10) for each pair of $x,\,y\in X$. ∎
Consider now the convex cone $\mathcal{C}_{+}$ of positive functions on $X$. Let $\alpha_{+}$ and $\beta_{+}$ be the corresponding functions associated to the convex cone $\mathcal{C}_{+}$ defined in (4.2). We denote by $\theta_{+}$ the projective metric associated to the convex cone $\mathcal{C}_{+}$ defined in (4.3). Now we establish the following result. Recall $\widehat{K}$ from (4.1).
@@
For all constants $\xi,\,a,\,v>0$, $1<b^{\prime}<b$, and $\lambda_{1}\in(0,1)$, we have $\theta_{+}(v_{1},v_{2})=\log(\sup\{v_{2}(x)/v_{1}(x):x\in X\})-\log(\inf\{v_{2}(y)/v_{1}(y):y\in X\})\leqslant\theta_{a,v,b}(v_{1},v_{2})$ for each pair of $v_{1},\,v_{2}\in\mathcal{C}(a,v,b)$, and
(4.12) $\operatorname{diam}_{\theta_{a,v,b}}(\mathcal{C}(\lambda_{1}a,v,b^{\prime}))\leqslant\widehat{K}(\lambda_{1},b,b^{\prime}).$
@@
Consider a pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(\lambda_{1}a,v,b^{\prime})\subseteq\mathcal{C}_{+}$. By the definition of $\mathcal{C}(\lambda_{1}a,v,b^{\prime})$, we have
$\varphi_{2}(y)\leqslant\varphi_{2}(x)\exp(a\lambda_{1}\rho(x,y)^{v})\quad\text{and}\quad\varphi_{1}(y)\geqslant\varphi_{1}(x)\exp(-a\lambda_{1}\rho(x,y)^{v})$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. Hence,
$\frac{\exp(a\rho(x,y)^{v})\varphi_{2}(x)-\varphi_{2}(y)}{\exp(a\rho(x,y)^{v})\varphi_{1}(x)-\varphi_{1}(y)}$ $\geqslant\frac{\varphi_{2}(x)}{\varphi_{1}(x)}\cdot\frac{\exp(a\rho(x,y)^{v})-\exp(a\lambda_{1}\rho(x,y)^{v})}{\exp(a\rho(x,y)^{v})-\exp(-a\lambda_{1}\rho(x,y)^{v})}$
(4.13) $\geqslant\inf\limits_{z>1}\left\{\frac{z-z^{\lambda_{1}}}{z-z^{-\lambda_{1}}}\right\}\cdot\frac{\varphi_{2}(x)}{\varphi_{1}(x)}$
$\geqslant\frac{1-\lambda_{1}}{1+\lambda_{1}}\cdot\frac{\varphi_{2}(x)}{\varphi_{1}(x)}$
for each pair of $x,\,y\in X$ with $\rho(x,y)\leqslant\xi$. By Definition 4.4, we have
$\varphi_{2}(y)\leqslant b^{\prime}\varphi_{2}(x)<b\varphi_{2}(x)\quad\text{and}\quad 0<\varphi_{1}(y)<b\varphi_{1}(x)$
@@
(4.14) $\frac{b\varphi_{2}(x)-\varphi_{2}(y)}{b\varphi_{1}(x)-\varphi_{1}(y)}\geqslant\frac{b-b^{\prime}}{b}\cdot\frac{\varphi_{2}(x)}{\varphi_{1}(x)}$
for each pair of $x,\,y\in X$. Consider an arbitrary pair of $v_{1},\,v_{2}\in\mathcal{C}(a,v,b)$. By the definition of $\alpha_{+}$ and $\beta_{+}$, we have $\alpha_{+}(v_{1},v_{2})=\inf_{t\in X}(v_{2}(t)/v_{1}(t))$, and $\beta_{+}(v_{1},v_{2})=\sup_{s\in X}(v_{2}(s)/v_{1}(s))$. Hence, $\theta_{+}(v_{1},v_{2})=\log(\sup\{v_{2}(x)/v_{1}(x):x\in X\})-\log(\inf\{v_{2}(y)/v_{1}(y):y\in X\}).$ By (4.4), we have $\alpha_{+}(v_{1},v_{2})\geqslant\alpha_{a,v,b}(v_{1},v_{2})$. Similarly, we have $\beta_{+}(v_{1},v_{2})\leqslant\beta_{a,v,b}(v_{1},v_{2})$. Therefore, $\theta_{+}(v_{1},v_{2})\leqslant\theta_{a,v,b}(v_{1},v_{2})$. Consider an arbitrary pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(\lambda_{1}a,v,b^{\prime})$. It follows from (4.4), (4.13), and (4.14) that $\alpha_{a,v,b}(\varphi_{1},\varphi_{2})\geqslant\frac{1-\lambda_{1}}{1+\lambda_{1}}\cdot\frac{b-b^{\prime}}{b}\cdot\alpha_{+}(\varphi_{1},\varphi_{2})$. Similarly, we can obtain that $\beta_{a,v,b}(\varphi_{1},\varphi_{2})\leqslant\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b}{b-b^{\prime}}\cdot\beta_{+}(\varphi_{1},\varphi_{2})$. Moreover, by the definitions of $\theta_{a,v,b}$ and $\theta_{+}$, it follows from (4.1) that
$\theta_{a,v,b}(\varphi_{1},\varphi_{2})$ $=\log(\beta_{a,v,b}(\varphi_{1},\varphi_{2}))-\log(\alpha_{a,v,b}(\varphi_{1},\varphi_{2}))$
$\leqslant\log(\beta_{+}(\varphi_{1},\varphi_{2}))-\log(\alpha_{+}(\varphi_{1},\varphi_{2}))+2\log\biggl{(}\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b}{b-b^{\prime}}\biggr{)}$
$=\theta_{+}(\varphi_{1},\varphi_{2})+\widehat{K}(\lambda_{1},b,b^{\prime})$
for each pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(\lambda_{1}a,v,b^{\prime})$. This completes the proof of (4.12). ∎
@@
[…]
$L^{m}(\mathcal{C}(a^{\prime},v_{0},C^{\prime}))\subseteq\mathcal{C}\biggl{(}\lambda_{1}a^{\prime},v_{0},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)},$
where $L$ is defined in (4.5). Thus, it follows from Proposition 4.6 that
(4.15) $\operatorname{diam}_{\theta_{a^{\prime},v_{0},C^{\prime}}}(L^{m}(\mathcal{C}(a^{\prime},v_{0},C^{\prime})))$ $\leqslant\operatorname{diam}_{\theta_{a^{\prime},v_{0},C^{\prime}}}\biggl{(}\mathcal{C}\biggl{(}\lambda_{1}a^{\prime},v_{0},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)}\biggr{)}$
$\leqslant\widehat{K}\biggl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)}.$
Obviously, we have $\mathbbm{1}\in\mathcal{C}(a^{\prime},v_{0},C^{\prime})$. By (2.16) and (2.17), it follows from $a^{\prime}>\frac{a_{0}}{\lambda^{v_{0}}-1}$ and $C^{2}<C^{\prime}$ that $1/u_{\varphi}\in\mathcal{C}(a^{\prime},v_{0},C^{\prime})$. Then by Proposition 4.3 for $L^{m}\colon\mathcal{C}(a^{\prime},v_{0},C^{\prime})\rightarrow\mathcal{C}(a^{\prime},v_{0},C^{\prime})$ and (4.15), we have
(4.16) $\theta_{a^{\prime},v_{0},C^{\prime}}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}$ $\leqslant(1-\exp(-Z))^{k-1}\cdot\theta_{a^{\prime},v_{0},C^{\prime}}\bigl{(}L^{m}(1/u_{\varphi}),\mathbbm{1}\bigr{)}$
$\leqslant(1-\exp(-Z))^{k-1}Z$
for each $k\in\mathbb{N}$, where $Z\coloneqq\widehat{K}\bigl{(}\lambda_{1},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\bigr{)}$. By Proposition 4.6, we have
(4.17) $\theta_{+}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}\leqslant\theta_{a^{\prime},v_{0},C^{\prime}}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}.$
@@
$\sup_{x\in X}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}=\sup_{x\in X}\{L^{mk}(1/{u_{\varphi}(x)})\}\geqslant 1\geqslant\inf_{x\in X}\{L^{mk}(1/{u_{\varphi}(x)})\}=\inf_{x\in X}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}.$
Hence, by Proposition 4.6, (2.17) and (4.18), we have
@@
for each $k\in\mathbb{N}$ satisfying $0\leqslant Z(1-\exp(-Z))^{k-1}\leqslant 1$. The proof is, therefore, complete. ∎

### v3

@@ 4.1. Ruelle–Perron–Frobenius operators and cones
Since $\varphi\in C^{0,v_{0}}(X,\rho)$, the function $\overline{\varphi}\in C^{0,v_{0}}(X,\rho)$ satisfies $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$. Hence, we have that
(4.8) $\exp(\overline{\varphi}(\overline{x}))\leqslant\exp(a_{0}\rho(\overline{x},y(\overline{x}))^{v_{0}})\cdot\exp(\overline{\varphi}(y(\overline{x})))\leqslant\exp(a_{0}\lambda^{-v_{0}}\rho(x,y)^{v_{0}})\cdot\exp(\overline{\varphi}(y(\overline{x})))$
for each $\overline{x}\in T^{-1}(x)$. Moreover, by (2.17), for each $\overline{x}\in T^{-1}(x)$,
@@
for each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$, where the last equality holds due to the definitions of $\lambda_{1}$ and $a^{\prime}$. Hence, $L(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$. By induction, it follows from $\lambda_{1}<1$ that for each $s\in\mathbb{N}$, $L^{s}(u)$ satisfies property (ii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$.
Finally, we establish that $L^{m}(u)$ satisfies property (iii) in Definition 4.4 for $\mathcal{C}(\lambda_{1}a^{\prime},b^{\prime})$, namely,
@@
Note that $L^{m}\colon C(X)\rightarrow C(X)$ is a linear operator, we can additionally assume that $\inf_{x\in X}u(x)=1$ without loss of generality. Since $u\in\mathcal{C}(a^{\prime},b)$, there exists a pair of $x_{*},\,x^{*}\in X$ with $u(x_{*})=1$ and $u(x^{*})=\sup_{x\in X}u(x)\leqslant b$. Depending on the value of $u(x^{*})$, we have the following two cases.
Case 1: $b^{\prime}\leqslant u(x^{*})\leqslant b$.
Consider a point $w\in X$, we shall show that $L^{m}(u)(w)\leqslant b^{\prime}$. Since $G^{\prime}$ is an $(\epsilon/2)$-net of $X$, there exists an integer $i\in[1,n^{\prime}]$ with $x_{*}\in B(x^{\prime}_{i},\epsilon/2)$. Since $T^{m}(B(x^{\prime}_{i},\epsilon/2))=X$, there exists $\overline{w}_{*}\in T^{-m}(w)\cap B(x_{i}^{\prime},\epsilon/2)$. Hence, $\rho(\overline{w}_{*},x_{*})\leqslant\rho(\overline{w}_{*},x^{\prime}_{i})+\rho(x^{\prime}_{i},x_{*})<\epsilon<\xi$. Moreover, since $u\in\mathcal{C}(a^{\prime},b)$, we have
@@
Hence, by the definitions of $\epsilon,\,Q,\,b^{\prime}$, $L(\mathbbm{1})=\mathbbm{1}$, (2.18), (4.5), and (4.11), we have
$L^{m}(u)(w)$ $=\frac{u(\overline{w}_{*})u_{\varphi}(\overline{w}_{*})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w}_{*})}+\sum_{\overline{w}\in T^{-m}(w)\smallsetminus\{\overline{w}_{*}\}}\frac{u(\overline{w})u_{\varphi}(\overline{w})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w})}$
$\leqslant\frac{\exp(a^{\prime}\epsilon^{v_{0}})\cdot u_{\varphi}(\overline{w}_{*})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w}_{*})}+u(x^{*})\sum_{\overline{w}\in T^{-m}(w)\smallsetminus\{\overline{w}_{*}\}}\frac{u_{\varphi}(\overline{w})}{u_{\varphi}(w)}\cdot e^{S_{m}\overline{\varphi}(\overline{w})}$
@@
$\leqslant b-(b^{\prime}-b/2)C^{-2}Q^{m}$
$=b^{\prime},$
where the last equality comes from a direct calculation.
Case 2: $u(x^{*})<b^{\prime}$.
Since $L(\mathbbm{1})=\mathbbm{1}$, we obtain that $\sup_{x\in X}L^{m}(u)(x)\leqslant\sup_{x\in X}u(x)<b^{\prime}.$
To summarize, we always have that $\sup_{x\in X}L^{m}(u)(x)\leqslant b^{\prime}$. Moreover, it follows from $L(\mathbbm{1})=\mathbbm{1}$ that $\inf_{x\in X}L^{m}(u)(x)\geqslant 1$. Therefore, we have estalished (4.10) for each pair of $x,\,y\in X$.
Therefore, we obtain (4.6) and complete the proof of Proposition 4.5. ∎
Let $\mathcal{C}_{+}$ be the set of positive functions on $X$. It is not hard to show that $\mathcal{C}_{+}$ is a convex cone. Let $\alpha_{+}$ and $\beta_{+}$ be the corresponding functions associated to $\mathcal{C}_{+}$ defined in (4.2), and $\theta_{+}$ be the projective metric associated to $\mathcal{C}_{+}$ defined in (4.3). Now, we establish the following result. Recall $\widehat{K}$ from (4.1).
@@
Fix constants $a>0$, $b>1$, and $0<\lambda_{1}<1$. Then
(4.12) $\theta_{+}(\varphi_{1},\varphi_{2})=\log(\sup\{\varphi_{2}(x)/\varphi_{1}(x):x\in X\})-\log(\inf\{\varphi_{2}(y)/\varphi_{1}(y):y\in X\})\leqslant\theta_{a,b}(\varphi_{1},\varphi_{2})$
for each pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(a,b)$, and
(4.13) $\operatorname{diam}_{\theta_{a,b}}(\mathcal{C}(\lambda_{1}a,b^{\prime}))\leqslant\widehat{K}(\lambda_{1},b,b^{\prime}).$
@@
Consider a pair of $\varphi_{1},\,\varphi_{2}\in\mathcal{C}(a,b)$. By the definition of $\alpha_{+}$ and $\beta_{+}$, we have that $\alpha_{+}(\varphi_{1},\varphi_{2})=\inf_{t\in X}(\varphi_{2}(t)/\varphi_{1}(t))$, and $\beta_{+}(\varphi_{1},\varphi_{2})=\sup_{s\in X}(\varphi_{2}(s)/\varphi_{1}(s))$. Hence,
$\theta_{+}(\varphi_{1},\varphi_{2})=\log(\sup\{\varphi_{2}(x)/\varphi_{1}(x):x\in X\})-\log(\inf\{\varphi_{2}(y)/\varphi_{1}(y):y\in X\}).$
By (4.4), we have $\alpha_{+}(\varphi_{1},\varphi_{2})\geqslant\alpha_{a,b}(\varphi_{1},\varphi_{2})$. Similarly, we have $\beta_{+}(\varphi_{1},\varphi_{2})\leqslant\beta_{a,b}(\varphi_{1},\varphi_{2})$. Hence, $\theta_{+}(\varphi_{1},\varphi_{2})\leqslant\theta_{a,b}(\varphi_{1},\varphi_{2})$.
Consider a pair of $\phi_{1},\,\phi_{2}\in\mathcal{C}(\lambda_{1}a,b^{\prime})\subseteq\mathcal{C}_{+}$. By the definition of $\mathcal{C}(\lambda_{1}a,b^{\prime})$, we have that
$\phi_{2}(y)\leqslant\phi_{2}(x)\exp(a\lambda_{1}\rho(x,y)^{v_{0}})\quad\text{ and }\quad\phi_{1}(y)\geqslant\phi_{1}(x)\exp(-a\lambda_{1}\rho(x,y)^{v_{0}})$
for each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$. Hence, we obtain that
$\frac{\exp(a\rho(x,y)^{v_{0}})\phi_{2}(x)-\phi_{2}(y)}{\exp(a\rho(x,y)^{v_{0}})\phi_{1}(x)-\phi_{1}(y)}$ $\geqslant\frac{\phi_{2}(x)}{\phi_{1}(x)}\cdot\frac{\exp(a\rho(x,y)^{v_{0}})-\exp(a\lambda_{1}\rho(x,y)^{v_{0}})}{\exp(a\rho(x,y)^{v_{0}})-\exp(-a\lambda_{1}\rho(x,y)^{v_{0}})}$
(4.14) $\geqslant\inf\limits_{z>1}\left\{\frac{z-z^{\lambda_{1}}}{z-z^{-\lambda_{1}}}\right\}\cdot\frac{\phi_{2}(x)}{\phi_{1}(x)}\geqslant\frac{1-\lambda_{1}}{1+\lambda_{1}}\cdot\frac{\phi_{2}(x)}{\phi_{1}(x)}$
for each pair of $x,\,y\in X$ with $\rho(x,y)<\xi$. By Definition 4.4, we have that
$\phi_{2}(y)\leqslant b^{\prime}\phi_{2}(x)<b\phi_{2}(x)\quad\text{ and }\quad 0<\phi_{1}(y)<b\phi_{1}(x)$
@@
(4.15) $\frac{b\phi_{2}(x)-\phi_{2}(y)}{b\phi_{1}(x)-\phi_{1}(y)}\geqslant\frac{b-b^{\prime}}{b}\cdot\frac{\phi_{2}(x)}{\phi_{1}(x)}$
for each pair of $x,\,y\in X$.
By (4.4), (4.14), and (4.15), we obtain that $\alpha_{a,b}(\phi_{1},\phi_{2})\geqslant\frac{1-\lambda_{1}}{1+\lambda_{1}}\cdot\frac{b-b^{\prime}}{b}\cdot\alpha_{+}(\phi_{1},\phi_{2})$. Similarly, we can obtain that $\beta_{a,b}(\phi_{1},\phi_{2})\leqslant\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b}{b-b^{\prime}}\cdot\beta_{+}(\phi_{1},\phi_{2})$. Moreover, since $\phi_{1},\,\phi_{2}\in\mathcal{C}(\lambda_{1}a,b^{\prime})$, we have $\frac{1}{b}\leqslant\frac{\phi_{2}(x)}{\phi_{1}(x)}\leqslant b$ for each $x\in X$. Hence, it follows from (4.1) that
$\theta_{a,b}(\phi_{1},\phi_{2})$ $=\log(\beta_{a,b}(\phi_{1},\phi_{2}))-\log(\alpha_{a,b}(\phi_{1},\phi_{2}))$
$\leqslant\log(\beta_{+}(\phi_{1},\phi_{2}))-\log(\alpha_{+}(\phi_{1},\phi_{2}))+2\log\biggl{(}\frac{1+\lambda_{1}}{1-\lambda_{1}}\cdot\frac{b}{b-b^{\prime}}\biggr{)}\leqslant\widehat{K}(\lambda_{1},b,b^{\prime})$
for each pair of $\phi_{1},\,\phi_{2}\in\mathcal{C}(\lambda_{1}a,b^{\prime})$. This completes the proof of (4.13). ∎
@@
Recall that $\lambda_{1}\coloneqq\frac{1+\lambda^{-v_{0}}}{2},\,a^{\prime}\coloneqq\frac{6a_{0}\lambda^{v_{0}}-2a_{0}}{(\lambda^{v_{0}}-1)^{2}}$, $C^{\prime}>\max\{C^{2},\,2\exp(a^{\prime})\}$, where $C$ is from Lemma 2.8. Write $D\coloneqq\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}$. Then by Proposition 4.5, we have $L^{m}(\mathcal{C}(a^{\prime},C^{\prime}))\subseteq\mathcal{C}(\lambda_{1}a^{\prime},D)$, where $L$ is defined in (4.5). Thus, it follows from Proposition 4.6 that
(4.16) $\operatorname{diam}_{\theta_{a^{\prime},C^{\prime}}}(L^{m}(\mathcal{C}(a^{\prime},C^{\prime})))\leqslant\operatorname{diam}_{\theta_{a^{\prime},C^{\prime}}}(\mathcal{C}(\lambda_{1}a^{\prime},D))\leqslant\widehat{K}(\lambda_{1},C^{\prime},D).$
Obviously, we have $\mathbbm{1}\in\mathcal{C}(a^{\prime},C^{\prime})$. By (2.17) and (2.18), it follows from $a^{\prime}>\frac{a_{0}}{\lambda^{v_{0}}-1}$ and $C^{\prime}>C^{2}$ that $1/u_{\varphi}\in\mathcal{C}(a^{\prime},C^{\prime})$. Then by Proposition 4.3 for $L^{m}\colon\mathcal{C}(a^{\prime},C^{\prime})\rightarrow\mathcal{C}(a^{\prime},C^{\prime})$ and (4.16), we have
(4.17) $\theta_{a^{\prime},C^{\prime}}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}\leqslant(1-\exp(-Z))^{k-1}\cdot\theta_{a^{\prime},C^{\prime}}(L^{m}(1/u_{\varphi}),\mathbbm{1})\leqslant(1-\exp(-Z))^{k-1}Z$
for each $k\in\mathbb{N}$, where $Z\coloneqq\widehat{K}(\lambda_{1},C^{\prime},D)$. By Proposition 4.6, we have
(4.18) $\theta_{+}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}\leqslant\theta_{a^{\prime},C^{\prime}}\bigl{(}L^{mk}(1/u_{\varphi}),\mathbbm{1}\bigr{)}.$
@@
$\sup_{x\in X}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}=\sup_{x\in X}\bigl{\{}L^{mk}(1/{u_{\varphi}(x)})\bigr{\}}\geqslant 1\geqslant\inf_{x\in X}\bigl{\{}L^{mk}(1/{u_{\varphi}(x)})\bigr{\}}=\inf_{x\in X}\frac{\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1})(x)}{u_{\varphi}(x)}.$
Hence, by (2.18), (4.12), and (4.19), we have
@@
$\leqslant C\bigl{(}\exp\bigl{(}\theta_{+}\bigl{(}\mathcal{L}^{mk}_{\overline{\varphi}}(\mathbbm{1}),u_{\varphi}\bigr{)}\bigr{)}-1\bigr{)}$
@@
for each $k\in\mathbb{N}$ with $0\leqslant Z(1-\exp(-Z))^{k-1}\leqslant 1$. Therefore, we complete the proof. ∎